# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 2.2 Empirical means of rescaled disjoint or sliding block maxima
for some remainder $R_{n}=o_{\mathbb{P}}(1)$. Loosely spoken, if the error term $R_{n}$ in the previous display is sufficiently small, the previous two displays and Theorem 2.3 imply the normal approximation
@@
Assembling terms and solving for $\hat{\theta}_{n}-\theta_{r}$, we obtain $\sqrt{\frac{n}{r}}a_{r}^{-2}(\hat{\theta}_{n}-\theta_{r})=\bar{\mathbb{G}}_{n,r}(h_{2}-(2Ph_{1})\cdot h_{1})+o_{\mathbb{P}}(1).$ The asymptotic variance of the limiting distribution has been explicitly calculated in Bücher and Staud, (2024), see their Corollary 4.1 and their Equations (C.1) and (C.2).
(ii) Probability weighted moments: consider the case $d=1$ and let ${\bm{\varphi}}_{n}:\mathbb{R}^{n}\to\mathbb{R}^{p}$ denote the vector containing the first $p$ empirical probability weighted moments. In that case, Equation (2.5) is met with $q=p$, ${\bm{\psi}}_{n}={\bm{\varphi}}_{n}$, $A(a,b)=\operatorname{diag}(a,\dots,a)$ and $B(a,b)=\operatorname{diag}(b,b/2,\dots,b/(p+1))$; see Formula (A.7) in Bücher and Zanger, (2023). Moreover, Equation (2.8) can be deduced from (A.5) in that paper (for $p=3$), with functions $h_{1}(x)=x$, $h_{2}(x)=xG_{\gamma}(x)+\int_{x}^{\infty}z\,\mathrm{d}G_{\gamma}(z)$ and $h_{3}(x)=xG_{\gamma}^{2}(x)+2\int_{x}^{\infty}zG_{\gamma}(z)\,\mathrm{d}G_{\gamma}(z)$. Finally, up to the treatment of a bias term, (2.10) corresponds to the assertion in their Theorem 3.2.
@@
In statistical applications, results like those in (2.9) are typically used as a starting point for inference, for instance in the form of confidence intervals for $\hat{\bm{\theta}}_{n}$. Routinely, such intervals would be based on normal approximations involving consistent estimators for the variances on the right-hand side of those displays, which in our case requires estimation of $\bm{a}_{r},\bm{b}_{r}$ and $\Sigma_{\bm{h}}=\Sigma_{\bm{h}}^{({\operatorname{mb}})}$. This can be a complicated task, especially for the sliding block maxima method in view of the complicated formula for the limiting variance. Alternatively, one can rely on bootstrap approximations instead of normal approximations, which can avoid the need for variance estimation. For the disjoint blocks case, the proof of Theorem 2.3 shows that the disjoint block maxima can essentially be considered independent (despite the serial dependence in the underlying observations). This suggests that the standard bootstrap based on resampling with replacement from the disjoint block maxima should be consistent; a conjecture that will be confirmed in Section 4. In the sliding block maxima case, however, the standard bootstrap cannot work since it destroys the serial dependence in the sliding block maxima sample. In classical time series analysis, a remedy consists of the block bootstrap, where blocks of successive observations are drawn with replacement. In our case, this heuristically requires the choice of a smoothing parameter sequence that is of larger order than the block size; see Remark 4.3 for details on the inconsistency of a naive approach to bootstrapping sliding block maxima. An alternative approach relies on the circular block maxima method as introduced in the next section.
@@ 3 The circular block maxima sample
In this section we introduce the circular block maxima method. We show that respective empirical means yield the same asymptotic variance as the sliding block maxima counterparts. The approach is interesting in its own (particularly for computational reasons), but most importantly for this paper it suggests a straightforward bootstrap approach for bootstrapping sliding block maxima that will be discussed in Section 4.
@@
$\mathcal{M}_{n,r}^{({\operatorname{cb}})}=(\bm{M}_{r,1}^{({\operatorname{cb}})},\dots,\bm{M}_{r,n}^{({\operatorname{cb}})})$
@@
where maxima over vectors in $\mathbb{R}^{d}$ are to be understood componentwise and where $[i:j]=\{i,i+1,\dots,j\}$. Next, for $\bm{U}=(\bm{u}_{1},\dots,\bm{u}_{kr})\in(\mathbb{R}^{d})^{kr}$, the circularization function is defined as
@@
It is constructive to take a closer look at the choices $k=1$ and $k=n/r$ (the number of disjoint blocks). In the former case, the circmax-sample is the same as the disjoint block maxima sample, but with every observation repeated exactly $r$ times. In the latter case, the first $n-r+1$ observations of the circmax-sample coincide with the sliding block maxima sample. New samples are obtained for every other choice of $k$, and in the subsequent developments we will mostly be interested in choices like $k=2$ or $k=3$.
@@
The following central result shows that the circmax-sample can be considered as a sample from the ‘correct’-limit distribution $G$ from (2.2) (a similar result on joint convergence of two circmax-observations can be found in Proposition 11.1). For $s\in\{1,\dots,n\}$, define
$\bm{Z}_{r,s}^{({\operatorname{cb}})}:=\frac{\bm{M}_{r,s}^{({\operatorname{cb}})}-\bm{b}_{r}}{\bm{a}_{r}}.$
@@
$\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\rightsquigarrow\bm{Z}\sim G\qquad(n\to\infty).$
Due to Proposition 3.1, statistical methods based on the circular block maxima method may be expected to work, asymptotically. As discussed in Section 2.2, a central ingredient for studying respective methods is weak convergence of empirical means. Adopting the notation from that section, we denote the empirical processes associated with the normalized sample $\bm{Z}_{r,1}^{({\operatorname{cb}})},\dots,\bm{Z}_{r,n}^{({\operatorname{cb}})}$ by
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P_{n,r}^{({\operatorname{cb}})}\big{)},\quad\tilde{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P_{r}\big{)},\quad\mathbb{G}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P\big{)},$ (3.1)
where $\mathbb{P}_{n,r}^{({\operatorname{cb}})}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{({\operatorname{cb}})}}$, $P_{n,r}^{({\operatorname{cb}})}=\frac{1}{kr}\sum_{s=1}^{kr}\mathbb{P}(\bm{Z}_{r,s}^{({\operatorname{cb}})}\in\cdot)$, $P_{r}=\mathbb{P}(\bm{Z}_{r,1}\in\cdot)$ and $P=\mathbb{P}(\bm{Z}\in\cdot)$. The following result can be regarded as a circmax-counterpart of Theorem 2.3.
@@
Suppose that Conditions 2.1 and 2.2 are met. Then, for fixed $k\in\mathbb{N}_{\geq 2}$ and any finite set of real valued functions $h_{1},\dots,h_{q}$ satisfying the integrability Condition 10.1(b) with $\nu>2/\omega$ with $\omega$ from Condition 2.2, we have
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}\bm{h}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{({\operatorname{sb}})})$
and $\lim_{n\to\infty}\operatorname{Cov}(\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}\bm{h})=\Sigma_{\bm{h}}^{({\operatorname{sb}})}$ with $\Sigma_{\bm{h}}^{({\operatorname{sb}})}$ from (2.4). Moreover, if the bias Condition 10.3 holds, then
$\tilde{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}\bm{h}\rightsquigarrow\mathcal{N}_{q}(\tfrac{D_{\bm{h},k}+E_{\bm{h}}}{k},\Sigma_{\bm{h}}^{({\operatorname{sb}})})$
(with $D_{\bm{h},k}$ often equal to zero, see Remark 3.3), and if additionally the bias Condition 10.2 is met, then
$\mathbb{G}_{n,r}^{({\operatorname{cb}})}\bm{h}\rightsquigarrow\mathcal{N}_{q}(B_{\bm{h}}+\tfrac{D_{\bm{h},k}+E_{\bm{h}}}{k},\Sigma_{\bm{h}}^{({\operatorname{sb}})}).$
Remark 3.3.
The bias term $D_{\bm{h},k}+E_{\bm{h}}$ is due to the fact that the rescaled circular block maxima $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\bm{X}_{1},\ldots,\bm{X}_{t})$ with index $s=(k-1)r+1+t$ for some $t\in\{1,\dots,r-1\}$ are not equal in distribution to a generic block maximum variable $\bm{M}_{r}$. This bias has two sources: $D_{\bm{h},k}$ is a measure for the average (over $t$) difference induced by the approximation of $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\bm{X}_{1},\ldots\bm{X}_{t})$ by $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{t})$ (with $(\tilde{\bm{X}}_{t})_{t}$ an independent copy of $(\tilde{\bm{X}}_{t})_{t}$), while $E_{\bm{h}}$ is a measure for the average difference induced by the approximation of $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{t})$ by that of a generic block maximum variable $\bm{M}_{r}$. In Lemma 10.4 we provide simple sufficient conditions for beta mixing time series that imply $D_{\bm{h},k}=B_{\bm{h}}=0$.
The proof of Theorem 3.2 (see in particular Proposition 11.1) shows that the sample of $kr$-blocks containing the circular block maxima can be considered asymptotically independent. This is akin to the sample of disjoint block maxima and, as discussed after Theorem 2.3, it suggests to define bootstrap versions of $\mathbb{G}_{n,r}^{({\operatorname{cb}})}$ by drawing $m(k)$-times from the sample of $kr$-blocks with replacement (which is computationally efficient: recall that all circular block maxima in a $kr$-block could be efficiently stored as a weighted sample with typically very few observations).
@@ 4 Bootstrapping block maxima estimators
Throughout this section, we discuss bootstrap approximations for the empirical processes from Section 3. For completeness, we also cover the disjoint block maxima case, as it can be identified with the circular block maxima sample for $k=1$.
For fixed $k\in\mathbb{N}$, consider the circmax-sample $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$. Independent of the observations, let $\bm{W}_{m(k)}=(W_{m(k),1},\dots,W_{m(k),m(k)})=(W_{1},\dots,W_{m(k)})$ be multinomially distributed with $m(k)$ trials and class probabilities $(m(k)^{-1},\dots,m(k)^{-1})$. Given the sample $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$, the bootstrap sample $\mathcal{M}_{n,r}^{({\operatorname{cb}}),*}$ is obtained by repeating the observations $(M_{r,s}^{({\operatorname{cb}})})_{s\in I_{kr,i}}$ from the $i$th $kr$-block exactly $W_{m(k),i}$-times, for every $i=1,\dots,m(k)$. The respective empirical measure of the rescaled bootstrap sample $\bm{Z}_{r,1}^{({\operatorname{cb}}),*},\dots,\bm{Z}_{r,n}^{({\operatorname{cb}}),*}$ with $\bm{Z}_{r,1}^{({\operatorname{cb}}),*}=(\bm{M}_{r,s}^{({\operatorname{cb}}),*}-\bm{b}_{r})/\bm{a}_{r}$ can then be written as
^$\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{({\operatorname{cb}}),*}}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}^{({\operatorname{cb}})}},$
@@
$\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}-\mathbb{P}_{n,r}^{({\operatorname{cb}})}).$ (4.1)
Note that the conditional distribution of $\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}$ given the data $\bm{X}_{1},\dots,\bm{X}_{n}$ can in practice be approximated to an arbitrary precision based on repeated sampling of $\bm{W}_{m(k)}$ (if $\bm{a}_{r},\bm{b}_{r}$ were known; we discuss extensions below). The following result shows that the bootstrap process $\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}$ provides a consistent distributional approximation for $\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}$ and for $\bar{\mathbb{G}}_{n,r}^{({\operatorname{sb}})}$.
@@
Suppose Conditions 2.1 and 2.2 are met. Then, for fixed $k\in\mathbb{N}_{\geq 2}$ and $\bm{h}$ satisfying the integrability Condition 10.1(b) with $\nu>2/\omega$ with $\omega$ from Condition 2.2, we have, for ${\operatorname{mb}}\in\{{\operatorname{sb}},{\operatorname{cb}}\}$,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1)$
as $n\to\infty$, where $d_{K}$ denotes the Kolmogorov metric for probability measures on $\mathbb{R}^{q}$. As a consequence, if the bias Condition 10.2 is met and if $r$ is chosen sufficiently large to guarantee that $B_{\bm{h}}=0$, we have
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\mathbb{G}_{n,r}^{({\operatorname{sb}})}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1),$
and if also the bias Condition 10.3 is met with $E_{\bm{h}}=D_{\bm{h},k}=0$, we have
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\mathbb{G}_{n,r}^{({\operatorname{cb}})}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1).$
Remarkably, the consistency of the circmax bootstrap for the sliding block maxima method does not require the bias Condition 10.3 (nor $E_{\bm{h}}=D_{\bm{h},k}=0$ in case it is met).
@@
Consider the disjoint block maxima sample, whose empirical means coincide with empirical means based on the circular block maxima sample with $k=1$. The results from Theorem 4.1 continue to hold for $k=1$ ($D_{\bm{h},k}=E_{\bm{h}}=0$ is then immediate), but with ‘${\operatorname{sb}}$’ replaced by ‘${\operatorname{db}}$’ at all instances.
@@
It seems natural to bootstrap sliding block maxima empirical means by the sliding block maxima analogue of (4.1), that is, by
$\hat{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{({\operatorname{sb}}),*}-P_{r}),\qquad\hat{\mathbb{P}}_{n,r}^{({\operatorname{sb}}),*}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{({\operatorname{sb}}),*}}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}^{({\operatorname{sb}})}}.$

### v2

@@ 2.2 Empirical means of rescaled disjoint or sliding block maxima
for some remainder $R_{n}=o_{\mathbb{P}}(1)$. Loosely spoken, if the error term $R_{n}$ is sufficiently small, the previous two displays and Theorem 2.3 imply the normal approximation
@@
Assembling terms and solving for $\hat{\theta}_{n}-\theta_{r}$, we obtain $\sqrt{\frac{n}{r}}a_{r}^{-2}(\hat{\theta}_{n}-\theta_{r})=\bar{\mathbb{G}}_{n,r}(h_{2}-(2Ph_{1})\cdot h_{1})+o_{\mathbb{P}}(1).$ The asymptotic variance of the limiting distribution has been explicitly calculated in BucSta24, see their Corollary 4.1 and Equations (C.1) and (C.2).
(ii) Probability weighted moments: consider the case $d=1$ and let ${\bm{\varphi}}_{n}:\mathbb{R}^{n}\to\mathbb{R}^{p}$ denote the vector containing the first $p$ empirical probability weighted moments. In that case, Equation (2.5) is met with $q=p$, ${\bm{\psi}}_{n}={\bm{\varphi}}_{n}$, $A(a,b)=\operatorname{diag}(a,\dots,a)$ and $B(a,b)=\operatorname{diag}(b,b/2,\dots,b/(p+1))$; see Formula (A.7) in BucZan23. Moreover, Equation (2.8) can be deduced from (A.5) in that paper (for $p=3$), with functions $h_{1}(x)=x$, $h_{2}(x)=xG_{\gamma}(x)+\int_{x}^{\infty}z\,\mathrm{d}G_{\gamma}(z)$ and $h_{3}(x)=xG_{\gamma}^{2}(x)+2\int_{x}^{\infty}zG_{\gamma}(z)\,\mathrm{d}G_{\gamma}(z)$. Finally, up to the treatment of a bias term, (2.10) corresponds to the assertion in their Theorem 3.2.
@@
In statistical applications, results like those in (2.9) are typically used as a starting point for inference, for instance in the form of confidence intervals for $\hat{\bm{\theta}}_{n}$. Routinely, such intervals would be based on normal approximations involving consistent estimators for the variances on the right-hand side of those displays, which in our case requires estimation of $\bm{a}_{r},\bm{b}_{r}$ and $\Sigma_{\bm{h}}=\Sigma_{\bm{h}}^{[{\operatorname{m}}]}$. This can be a complicated task, especially for the sliding block maxima method in view of the complicated formula for the limiting variance. Alternatively, one can rely on bootstrap approximations instead of normal approximations. However, as already illustrated in the introduction, a naive approach for bootstrapping sliding block maxima fails (see also Remark 4.3 below), which prompts us to first introduce the circular block maxima method in the next section.
@@ 3 The circular block maxima sample
In this section we introduce the circular block maxima method. We show that respective empirical means yield the same asymptotic variance as the sliding block maxima counterparts. The approach is interesting in its own right (particularly for computational reasons), but most importantly for this paper it suggests a straightforward bootstrap approach for bootstrapping sliding block maxima that will be discussed in Section 4.
@@
$\mathcal{M}_{n,r}^{[{\operatorname{c}}]}=(\bm{M}_{r,1}^{[{\operatorname{c}}]},\dots,\bm{M}_{r,n}^{[{\operatorname{c}}]})$
@@
where maxima over vectors in $\mathbb{R}^{d}$ are understood componentwise and where $[i:j]=\{i,i+1,\dots,j\}$. Next, for $\bm{U}=(\bm{u}_{1},\dots,\bm{u}_{kr})\in(\mathbb{R}^{d})^{kr}$, the circularization function is defined as
@@
The entire construction is also summarized in pseudo-code in Algorithm 1.
Algorithm 1 (Univariate) circ-max Method
1:Sample $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$, block size $r$, blocking parameter $k$
2:function Circmax($x,r,k$)
3:     Partition: Divide $x$ into $m(k)=n/(kr)$ disjoint blocks of size $kr$.
4:     for each block $(x_{(i-1)kr+1},\dots,x_{ikr})\in\mathbb{R}^{kr}$ with $i=1,\dots,m(k)$ do
5:         $\mathcal{X}_{kr,i}^{\text{circ}}\leftarrow(x_{(i-1)kr+1},\dots,x_{ikr},x_{(i-1)kr+1},\dots,x_{(i-1)kr+r-1})\in\mathbb{R}^{kr+r-1}$
6$\triangleright$ Circular trafo (repeat first $r-1$ observations)
7:         $(M_{r,s}^{[{\operatorname{c}}]})_{s\in I_{kr,i}}\leftarrow\text{slid-max}(\mathcal{X}_{kr,i}^{\text{circ}}\mid r)\in\mathbb{R}^{kr}$triangleright Sliding maxima of circular blocks
8:     end for
9:     $\mathcal{M}_{n,r}^{[{\operatorname{c}}]}\leftarrow(M_{r,1}^{[{\operatorname{c}}]},\dots,M_{r,n}^{[{\operatorname{c}}]})\in\mathbb{R}^{n}$triangleright Concatenate
10:     return $\mathcal{M}_{n,r}^{[{\operatorname{c}}]}$
11:end function
It is constructive to take a closer look at the choices $k=1$ and $k=n/r$ (the number of disjoint blocks). In the former case, the circmax-sample is the same as the disjoint block maxima sample, but with every observation repeated exactly $r$ times. In the latter case, the first $n-r+1$ observations of the circmax-sample coincide with the sliding block maxima sample, so the two are asymptotically equivalent. New samples are obtained for every other choice of $k$, and in the subsequent developments we will mostly be interested in choices like $k=2$ or $k=3$.
@@
The following central result shows that the circmax-sample can be considered as a sample from the ‘correct’ limit distribution $G$ from (2.2) (a similar result on joint convergence of two circmax-observations can be found in Proposition C.1 of the supplement). Let
$\bm{Z}_{r,s}^{[{\operatorname{c}}]}:=\frac{\bm{M}_{r,s}^{[{\operatorname{c}}]}-\bm{b}_{r}}{\bm{a}_{r}},\qquad s\in\{1,\dots,n\};$
@@
$\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{[{\operatorname{c}}]}\rightsquigarrow\bm{Z}\sim G\qquad(n\to\infty).$
Due to Proposition 3.1, statistical methods based on the circular block maxima method may be expected to work asymptotically. As discussed in Section 2.2, a central ingredient for studying respective methods is weak convergence of empirical means. Adopting the notation from that section, we denote the empirical processes associated with the normalized sample $\bm{Z}_{r,1}^{[{\operatorname{c}}]},\dots,\bm{Z}_{r,n}^{[{\operatorname{c}}]}$ by
$\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{[{\operatorname{c}}]}-P_{n,r}^{[{\operatorname{c}}]}\big{)},\quad\tilde{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{[{\operatorname{c}}]}-P_{r}\big{)},\quad\mathbb{G}_{n,r}^{[{\operatorname{c}}]}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{[{\operatorname{c}}]}-P\big{)},$ (3.1)
where $\mathbb{P}_{n,r}^{[{\operatorname{c}}]}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{[{\operatorname{c}}]}}$, $P_{n,r}^{[{\operatorname{c}}]}=\frac{1}{kr}\sum_{s=1}^{kr}\mathbb{P}(\bm{Z}_{r,s}^{[{\operatorname{c}}]}\in\cdot)$, $P_{r}=\mathbb{P}(\bm{Z}_{r,1}\in\cdot)$ and $P=\mathbb{P}(\bm{Z}\in\cdot)$. The following result can be regarded as a circmax-counterpart of Theorem 2.3. Recall that the integrability and bias Conditions 9.1-9.3 have been postponed to Section 9 in the appendix, where we also provide an extended discussion of the bias in 9.4.
@@
Suppose that Conditions 2.1 and 2.2 are met. Then, for fixed $k\in\mathbb{N}_{\geq 2}$ and any finite set of real valued functions $h_{1},\dots,h_{q}$ satisfying the integrability Condition 9.1(b) with $\nu>2/\omega$, where $\omega$ is from Condition 2.2, we have $\lim_{n\to\infty}\operatorname{Cov}(\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}\bm{h})=\Sigma_{\bm{h}}^{[{\operatorname{s}}]}$ and
$\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}\bm{h}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{[{\operatorname{s}}]})$
with $\Sigma_{\bm{h}}^{[{\operatorname{s}}]}$ from (2.4). Moreover, if the bias Condition 9.3 holds, then
$\tilde{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}\bm{h}\rightsquigarrow\mathcal{N}_{q}(\tfrac{D_{\bm{h},k}+E_{\bm{h}}}{k},\Sigma_{\bm{h}}^{[{\operatorname{s}}]})$
(with $D_{\bm{h},k}$ often equal to zero, see Remark 9.4), and if additionally the bias Condition 9.2 is met, then
$\mathbb{G}_{n,r}^{[{\operatorname{c}}]}\bm{h}\rightsquigarrow\mathcal{N}_{q}(B_{\bm{h}}+\tfrac{D_{\bm{h},k}+E_{\bm{h}}}{k},\Sigma_{\bm{h}}^{[{\operatorname{s}}]}).$
The proof of Theorem 3.2 (see in particular Proposition C.1 in the supplement) shows that the sample of $kr$-blocks containing the circular block maxima can be considered asymptotically independent. This is akin to the sample of disjoint block maxima and, as discussed after Theorem 2.3, it suggests to define bootstrap versions of $\mathbb{G}_{n,r}^{[{\operatorname{c}}]}$ by drawing $m(k)$-times from the sample of $kr$-blocks with replacement (which is computationally efficient: recall that all circular block maxima in a $kr$-block could be efficiently stored as a weighted sample with typically very few observations).
@@ 4 Bootstrapping block maxima estimators
Throughout this section, we discuss bootstrap approximations for the empirical processes from Section 3. For completeness, we also cover the disjoint block maxima case, as it corresponds to the circular block maxima sample when $k=1$.
For fixed $k\in\mathbb{N}$, consider the circmax-sample $\mathcal{M}_{n,r}^{[{\operatorname{c}}]}$. Independent of the observations, let $\bm{W}_{m(k)}=(W_{m(k),1},\dots,W_{m(k),m(k)})=(W_{1},\dots,W_{m(k)})$ be multinomially distributed with $m(k)$ trials and class probabilities $(m(k)^{-1},\dots,m(k)^{-1})$. Given the sample $\mathcal{M}_{n,r}^{[{\operatorname{c}}]}$, the bootstrap sample $\mathcal{M}_{n,r}^{[{\operatorname{c}}],*}$ is obtained by repeating the observations $(M_{r,s}^{[{\operatorname{c}}]})_{s\in I_{kr,i}}$ from the $i$th $kr$-block exactly $W_{m(k),i}$-times, for every $i=1,\dots,m(k)$. The respective empirical measure of the rescaled bootstrap sample $\bm{Z}_{r,1}^{[{\operatorname{c}}],*},\dots,\bm{Z}_{r,n}^{[{\operatorname{c}}],*}$ with $\bm{Z}_{r,1}^{[{\operatorname{c}}],*}=(\bm{M}_{r,s}^{[{\operatorname{c}}],*}-\bm{b}_{r})/\bm{a}_{r}$ can then be written as
^$\hat{\mathbb{P}}_{n,r}^{[{\operatorname{c}}],*}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{[{\operatorname{c}}],*}}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}^{[{\operatorname{c}}]}},$
@@
$\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{[{\operatorname{c}}],*}-\mathbb{P}_{n,r}^{[{\operatorname{c}}]}).$ (4.1)
Note that the conditional distribution of $\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}$ given the data $\bm{X}_{1},\dots,\bm{X}_{n}$ can in practice be approximated to an arbitrary precision based on repeated sampling of $\bm{W}_{m(k)}$ (if $\bm{a}_{r},\bm{b}_{r}$ were known; we discuss extensions below). The following result shows that the bootstrap process $\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}$ provides a consistent distributional approximation for $\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}$ and for $\bar{\mathbb{G}}_{n,r}^{[{\operatorname{s}}]}$.
Again, note that the integrability and bias conditions 9.1-9.3 referenced in the results of this section can be found in the Section 9.
@@
Suppose Conditions 2.1 and 2.2 are met. Then, for fixed $k\in\mathbb{N}_{\geq 2}$ and $\bm{h}$ satisfying the integrability Condition 9.1(b) with $\nu>2/\omega$ where $\omega$ is from Condition 2.2, we have, for ${\operatorname{m}}\in\{{\operatorname{s}},{\operatorname{c}}\}$,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1)$
as $n\to\infty$, where $d_{K}$ denotes the Kolmogorov metric for probability measures on $\mathbb{R}^{q}$. As a consequence, if the bias Condition 9.2 is met and if $r$ is chosen sufficiently large to guarantee that $B_{\bm{h}}=0$, we have
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\mathbb{G}_{n,r}^{[{\operatorname{s}}]}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1),$
and if also the bias Condition 9.3 is met with $E_{\bm{h}}=D_{\bm{h},k}=0$, we have
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\mathbb{G}_{n,r}^{[{\operatorname{c}}]}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1).$
Remarkably, the consistency of the circmax bootstrap for the sliding block maxima method does not require the bias Condition 9.3 (nor $E_{\bm{h}}=D_{\bm{h},k}=0$ in case it is met).
@@
Consider the disjoint block maxima sample, whose empirical means coincide with empirical means based on the circular block maxima sample with $k=1$. The results from Theorem 4.1 continue to hold for $k=1$ ($D_{\bm{h},k}=E_{\bm{h}}=0$ is then immediate), but with ‘${\operatorname{s}}$’ replaced by ‘${\operatorname{d}}$’ at all instances.
@@
It seems natural to bootstrap sliding block maxima empirical means by blockwise resampling, that is, by the sliding block maxima analogue of (4.1),
$\hat{\mathbb{G}}_{n,r}^{[{\operatorname{s}}],*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{[{\operatorname{s}}],*}-P_{r}),\qquad\hat{\mathbb{P}}_{n,r}^{[{\operatorname{s}}],*}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{[{\operatorname{s}}],*}}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}^{[{\operatorname{s}}]}}.$